Hi,I would like to know the easiest solution to this question:
What is the size of an angle whose complement is 25% of its supplement
Thank You.

Answer by mananth(16946)

What is the size of an angle whose complement is 25% of its supplement
let angle be x
complement of x= (90-x)
Supplement = (180-x)
0.25(180-x)
90-x + 0.25(180-x)
90-x= 45-0.25x
90-45=x-0.25x
45=0.75x
45/0.75 =x
60=x
Check
complement = 90-60=30
0.25(180-60)
0.25(120)
30